In a group of quarters and nickels, there are four more nickels than quarters. How many nickels and quarters are there if the coins are worth $2.30
Answer by RAY100(1637)

Let n = # nickles,,,,,,and q= # quarters
.
n= q+4,,,,,,,,,given
.
(.05)n + (.25)q = 2.30,,,,,,,given
.
subst,,,,(.05)(q+4) + (.25)q = 2.30
.
.05q + .20 +.25q =2.30
.
.30q = 2.10
.
q=7,,,,,,n=7+4 =11
.
.
check,,, .05 *11 + .25*7 = .55+1.75 = 2.30,,,,,,,ok
.